# Diffusion Beats Autoregressive in Data-Constrained Settings

### v5

@@ Abstract
Autoregressive (AR) models have long dominated the landscape of large language models, driving progress across a wide range of tasks. Recently, diffusion-based language models have emerged as a promising alternative, though their advantages over AR models remain underexplored. In this paper, we systematically study masked diffusion models in data-constrained settings—where training involves repeated passes over limited data—and find that they significantly outperform AR models when compute is abundant but data is scarce. Diffusion models make better use of repeated data, achieving lower validation loss and superior downstream performance. We interpret this advantage as implicit data augmentation: masked diffusion exposes the model to a diverse distribution of token orderings and prediction tasks, unlike AR’s fixed left-to-right factorization. We find new scaling laws for diffusion models and derive a closed-form expression for the critical compute threshold at which diffusion begins to outperform AR. These results suggest that when data, not compute, is the bottleneck, diffusion models offer a compelling alternative to the standard AR paradigm. Our code is available at: https://diffusion-scaling.github.io.
@@ 4 Experiments
Our goal is to compare the performance of masked diffusion models and autoregressive models in data-constrained settings. To this end, we train a total of 200 models—100 diffusion models and 100 autoregressive models—across varying unique data sizes, model scales, and epoch counts. We present the empirical results in Section 4.1. In Section 4.2, we fit scaling laws tailored to data-constrained regimes for both model types, following the methodology introduced by Muennighoff et al.muennighoff2023scaling . These scaling laws allow us to analyze performance trends and identify scenarios where diffusion models should be preferred over autoregressive ones (Section 4.3). Finally, in Section 4.4, we demonstrate that the superior validation loss of diffusion models indeed correlates with improved downstream task performance.
@@ 4.5 Why do Diffusion models outperform AR models in data-constrained settings?
To test whether this diversity could be replicated in AR models through explicit augmentation, we first experimented with standard perturbation-based techniques during AR training. Specifically, we applied (i) attention dropout, where a random fraction of attention weights was dropped at varying levels (25%, 50%, 75%), and (ii) token masking, where a subset of input tokens was masked by zeroing out their attention weights across all layers, while retaining the next-token prediction objective. As shown in Figure 7(a) and Figure 7(b), neither of these augmentations improved validation loss. In all cases, AR models continued to overfit and remained significantly behind diffusion models trained for longer epochs.
We next investigated whether the key factor behind diffusion’s performance lies in its exposure to diverse token orderings. To evaluate this, we trained AR models with varying numbers of token orderings: $N=1$ corresponds to standard left-to-right training, while $N=k$ includes the left-to-right order along with k1k{-}1 additional random permutations of the token sequence. As shown in Figure 8, increasing $N$ consistently reduced validation loss and delayed overfitting. At $N=16$, the AR models approached the performance of diffusion models, suggesting that training over diverse orderings is indeed a major driver of diffusion’s data efficiency.
These findings support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained over a rich distribution of conditional prediction tasks. While AR models can partially recover this behavior via explicit random orderings, they remain fundamentally constrained by their fixed prediction structure.
Finally, this analysis suggests a natural interpolation between the two paradigms. By controlling the diversity of prediction tasks—through masking or reordering—it may be possible to design hybrid models that trade off between compute efficiency (AR-like behavior) and data efficiency (diffusion-like behavior). Exploring this continuum offers a promising direction for future research. More details on how we permute the sequence can be found in 11

### v6

@@ Abstract
Autoregressive (AR) models have long dominated the landscape of large language models, driving progress across a wide range of tasks. Recently, diffusion-based language models have emerged as a promising alternative, though their advantages over AR models remain underexplored. In this paper, we systematically study masked diffusion models in data-constrained settings—where training involves repeated passes over limited data—and find that they significantly outperform AR models when compute is abundant but data is scarce. Diffusion models make better use of repeated data, achieving lower validation loss and superior downstream performance. We find new scaling laws for diffusion models and derive a closed-form expression for the critical compute threshold at which diffusion begins to outperform AR. Finally, we explain why diffusion models excel in this regime: their randomized masking objective implicitly trains over a rich distribution of token orderings, acting as an implicit data augmentation that AR’s fixed left-to-right factorization lacks. Our results suggest that when data, not compute, is the bottleneck, diffusion models offer a compelling alternative to the standard AR paradigm. Our code is available at: https://diffusion-scaling.github.io.
@@ 1 Introduction
  6. 6.
    Exposure to different token orderings helps explain diffusion’s data efficiency. By adding explicit data augmentations to AR training, we find that diffusion models’ advantage arises from their exposure to a diverse set of token orderings. Essentially, the randomized masking in diffusion’s objective serves as implicit data augmentation, allowing it to generalize beyond the fixed left-to-right factorization of AR models. (Section 4.5)
@@ 4 Experiments
Our goal is to compare the performance of masked diffusion models and autoregressive models in data-constrained settings. To this end, we train a total of 200 models—100 diffusion models and 100 autoregressive models—across varying unique data sizes, model scales, and epoch counts. We present the empirical results in Section 4.1. In Section 4.2, we fit scaling laws tailored to data-constrained regimes for both model types, following the methodology introduced by Muennighoff et al.muennighoff2023scaling . These scaling laws allow us to analyze performance trends and identify scenarios where diffusion models should be preferred over autoregressive ones (Section 4.3). In Section 4.4, we demonstrate that the superior validation loss of diffusion models indeed correlates with improved downstream task performance. Finally, in Section 4.5 we investigate the underlying cause of diffusion’s advantage in data-constrained settings, showing that its exposure to diverse token orderings enables better generalization than AR’s fixed left-to-right factorization.
@@ 4.5 Why do Diffusion models outperform AR models in data-constrained settings?
To test whether the diversity in diffusion training could be replicated in AR models via explicit augmentation, we first applied standard perturbation-based techniques during AR training. Specifically, we used: (i) attention dropout — randomly dropping 25%, 50%, or 75% of attention weights; and (ii) token masking — masking a subset of input tokens by zeroing their attention weights across all layers, while retaining the standard next-token prediction objective.
As shown in Figures 7(a) and 7(b), neither approach improved validation loss. In all cases, AR models continued to overfit and remained far behind diffusion models trained for longer epochs. All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 4(a), trained for 500 epochs.
We next investigated whether diffusion’s advantage stems from exposure to diverse token orderings. To test this, we trained AR models with varying numbers of orderings: $N=1$ denotes standard left-to-right training, while $N=k$ adds k1k{-}1 random permutations of the sequence order. All the AR models in this setting had 278M parameters and were trained for 100 epochs.
As shown in Figure 8, increasing $N$ consistently lowered validation loss and delayed overfitting. At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s data efficiency.
These results support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained on a richer distribution of conditional prediction tasks.
Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and data efficiency (diffusion-like). Exploring this continuum is a promising direction for future work. Details of our permutation process are in Section 11.